The Uruguayan supply and demand for wheat are depicted below. The world price of wheat is $3 per bag, and Uruguay currently has a 100 percent tariff on wheat imports, so that its domestic wheat prices are double world prices. 3. 12 I3 100 4 0 2 345 678 9 1011 12 Q Suppose Uruguay forms a free trade area with Argentina, where the price of wheat is $5. Assuming that Uruguay is too small to influence the price of wheat in Argentina, complete the table below. Use minus signs to indicate negative changes Change in price of wheat faced by Uruguayan producers and consumers Change in Uruguayan tariff revenue Change in quantity of Uruguayan wheat imports Change in Uruguayan consumer surplus Change in Uruguayan produer surplus Net change in Uruguayan welfare
Answer #1

In following graph, world price (Pw) is $3, at which domestic demand is Q1 (= 11), domestic supply is Q2 (= 1) and imports equal (Q1 - Q2) (= 11 - 1 = 10). Post-tariff price (Pt) is $6, at which domestic demand is Q3 (= 8), domestic supply is Q4 (= 4) and imports equal (Q3 - Q4) (= 8 - 4 = 4).

1 1183 2 21098 11198 76543210

(i) Change in price = Pt - Pw = $6 - $3 = $3 (Increase)

(ii) Change in tariff revenue = Area CDFE = $3 x (8 - 4) = $3 x 4 = $12 (Increase)

(iii) Change in imports = 10 - 4 = 6 (Decrease)

(iv) Change in consumer surplus (CS) = Area PwACPt = (1/2) x $3 x (11 + 8)** = $1.5 x 19 = $28.5 (Decrease)

(v) Change in producer surplus (PS) = Area PwBDPt = (1/2) x $3 x (1 + 4)** = $1.5 x 5 = $7.5 (Increase)

(vi) Net change in welfare ($) = Decrease in CS - Increase in PS - Tariff revenue = 28.5 - 12 - 7.5 = 9 (Decrease)

**Area of trapezium = (1/2) x Height s Sum of parallel sides